# Asymmetric thin-shell wormholes in the Kalb-Ramond background: Observational characteristics and extra photon rings

### v2

@@ IV Conclusion and Discussion
This paper primarily investigates the observational characteristics of ATSW and BH under the same mass parameter $M$, charge $Q$, and Lorentz-violating parameter $l$ in the KR field. Unlike BH, ATSW connects two spacetimes with mass parameters $M_{1}$ and $M_{2}$ through a throat, assuming that the observer at infinity is located in the spacetime with mass parameter $M_{1}$. Initially, we calculate the variations in the photon sphere radius and critical impact parameter $b_{c}$ in the ATSW spacetime under different values of charge $Q$ and Lorentz-violating parameter $l$, as shown in Table 1 and Table 2. The results indicate that as the charge $Q$ and Lorentz-violating parameter $l$ increase, the photon sphere radius $r_{p}$ and critical impact parameter $b_{c}$ decrease accordingly. Next, we plot the relationship between the effective potential energy of ATSW and BH and the variations in charge $Q$ and Lorentz-violating parameter $l$, as illustrated in Fig.1 and Fig.2. The photon trajectories can be categorized into three scenarios: when $b_{1}<Zb_{c_{2}}$, photons fall from the M1 spacetime into the M2 spacetime and move to infinity; when $Zb_{c_{2}}<b_{1}<b_{c_{1}}$, photons enter the potential barrier of the M2 spacetime from the M1 spacetime and then return to the $M_{1}$ spacetime; when $b_{1}>b_{c_{1}}$, photons reach the potential barrier of the $M_{1}$ spacetime and then move towards infinity in the $M_{1}$ spacetime.
To better understand the photon trajectories and deflection angles, we have plotted the photon trajectories and deflection angles under different values of charge $Q$ and Lorentz-violating parameter $l$, as shown in Fig.3 and Fig.4. The results indicate that $b_{c_{1}}$ decreases with the increase of $Q$ and $l$, whereas $Zb_{c_{2}}$ exhibits an opposite trend. It is also observed that the photon trajectories are significantly influenced by the Lorentz-violating parameter $l$. Additionally, we have calculated the transfer functions of ATSW and found that, unlike BH, ATSW possesses new second and third transfer functions, which correspond to the lens ring and photon ring group, respectively. The unique significance of the new transfer functions in the ATSW spacetime is that when $Zb_{c_{2}}<b_{1}<b_{c_{1}}$, the spacetime $M_{2}$ can return photons to the spacetime $M_{1}$.

### v3

@@ IV Conclusion and Discussion
This paper primarily investigates the observational characteristics of the ATSW and BH under the same mass parameter $M$, charge $Q$, and Lorentz-violation parameter $l$ in the KR field. Unlike BH, the ATSW connects two spacetimes with mass parameters $M_{1}$ and $M_{2}$ through a throat, assuming that the observer at infinity is located in the spacetime with mass parameter $M_{1}$. Initially, we calculate the variations in the photon sphere radius and critical impact parameter $b_{c}$ in the ATSW spacetime under different values of charge $Q$ and Lorentz-violation parameter $l$, as shown in Table. (1) and Table. (2). The results indicate that as the charge $Q$ and Lorentz-violation parameter $l$ increase, the photon sphere radius $r_{p}$ and critical impact parameter $b_{c}$ decrease accordingly. Next, we plot the relationship between the effective potential energy of the ATSW and BH and the variations in charge $Q$ and Lorentz-violation parameter $l$, as illustrated in Fig.3 and 4. The photon trajectories can be categorized into three scenarios: when $b_{1}<Zb_{c_{2}}$, photons fall from the $M_{1}$ spacetime into the $M_{2}$ spacetime and move to infinity; when $Zb_{c_{2}}<b_{1}<b_{c_{1}}$, photons enter the potential barrier of the $M_{2}$ spacetime from the $M_{1}$ spacetime and then return to the $M_{1}$ spacetime; when $b_{1}>b_{c_{1}}$, photons reach the potential barrier of the $M_{1}$ spacetime and then move towards infinity in the $M_{1}$ spacetime.
To better understand the photon trajectories and deflection angles, we have plotted the photon trajectories and deflection angles under different values of charge $Q$ and Lorentz-violation parameter $l$, as shown in Figs.5 and 6. The results indicate that $b_{c_{1}}$ decreases with the increase of $Q$ and $l$, whereas $Zb_{c_{2}}$ exhibits an opposite trend. It is also observed that the photon trajectories are significantly influenced by the Lorentz-violation parameter $l$. Additionally, we have calculated the transfer functions of ATSW and found that, unlike BH, ATSW possesses new second and third transfer functions, which correspond to the lens ring and photon ring group, respectively. The unique significance of the new transfer functions in the ATSW spacetime is that when $Zb_{c_{2}}<b_{1}<b_{c_{1}}$, the spacetime $M_{2}$ can return photons to the spacetime $M_{1}$.